OpenStudy (anonymous):

The difference between the lengths of the major axis and the latus rectum of an ellipse is option are a.(2,1) b.(3,4) c.(-4,-3) d.(5,6)

OpenStudy (agent0smith):

Was there more to this question? major axis length = 2a latus rectum = 2b^2/a Difference between the lengths is just one subtract the other: \[L = 2a - \frac{ 2b^2 }{ a } \]which you can simplify in a few ways... \[L = 2a - \frac{ 2b^2 }{ a } = 2\frac{ a^2 - b^2 }{a }=2\frac{ c^2 }{ a }\]

OpenStudy (anonymous):

so the answer is not in the option i dont see a near 1

OpenStudy (agent0smith):

Well the answers don't seem to make sense for the question... all the answers given are co-ordinates of points (x, y), points are not lengths.